# CSI-BERT2: A BERT-inspired Framework for Efficient CSI Prediction and Classification in Wireless Communication and Sensing

### v3

@@ IV-C2 CSI Prediction Task
Given the scarcity of real CSI datasets in wireless communications, we adopt the widely used approach of leveraging wireless sensing datasets for performance evaluation. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [53], RNN [55], GRU [57], and Mamba [58], which are also widely used in CSI prediction [12, 13, 14, 15, 59, 60]. We also compare our model with some CSI prediction networks like OCEAN [61], CV-3DCNN [62], and ConvLSTM [63]. The prediction error and time cost (using GPU) are shown in Table VI. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its prediction time for the entire testing set across different datasets remains under 0.5 seconds, which is acceptable considering that the testing set takes several minutes. Additionally, we provide the Floating Point Operations (FLOPs) of different methods for a single input in the table. Our CSI-BERT2 model requires higher FLOPs, but it only reaches around 0.5G, making it suitable for real-time applications in edge devices [64, 65].
@@
Configuration CommPre WiGesture [11]
@@
LSTM [53, 12, 13, 14, 15] 133K 0.0134 5.1107 0.4849 0.5722 0.03 12.3254 0.2397 0.3967 0.05
RNN [55, 59] 33K 0.0033 4.6164 0.4607 0.5385 0.03 19.4708 0.2877 0.4063 0.04
GRU [57, 60] 100K 0.0100 5.4009 0.5128 0.6367 0.04 19.7180 0.2922 0.4243 0.04
Mamba [58] 5M 0.5062 5.1286 0.4430 0.4838 0.23 12.3281 0.2392 0.3277 0.24
OCEAN [61] 126K 0.0126 5.6013 0.4863 0.5576 0.05 19.6257 0.2925 0.4231 0.05
CV-3DCNN [62] 19K 0.1033 4.5276 0.4794 0.6810 0.04 11.3017 0.2267 0.3044 0.04
ConvLSTM [63] 152K 0.1136 5.0015 0.4898 0.6031 0.04 19.7038 0.2921 0.4242 0.04
Configuration WiFall[41] WiCount
@@
LSTM [53, 12, 13, 14, 15] 133K 0.0134 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05
RNN [55, 59] 33K 0.0033 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03
GRU [57, 60] 100K 0.0100 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02
Mamba [58] 5M 0.5062 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11
OCEAN [61] 126K 0.0126 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02
CV-3DCNN [62] 19K 0.1033 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02
ConvLSTM [63] 152K 0.1136 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02
@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table V, where it achieves the best performance across all scenarios. Notably, the original CSI-BERT fails to outperform models such as Chen et al. [56] and WiGrumt [54] in many scenarios. However, our CSI-BERT2 shows substantial improvement due to the enhanced model structure, thanks to the ARL capturing the relationships and importance among subcarriers, and the MLP-based time embedding layer that encodes time information more effectively. Moreover, aided by the time embedding mechanism, CSI-BERT2 can be effectively applied in scenarios where the training data includes samples with varying sampling rates, or where the sampling rate of the testing set differs from that of the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table VII, we compare the performance of different models where the training and testing sets both have 100Hz and 50Hz CSI, as well as where the training set only has 100Hz CSI but the testing set has 50Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing to function properly in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of time embedding.
@@
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% Person Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [55] LSTM [53] Chen et al. [56] WiGRUNT [54] CSI-BERT [11] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% Person Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D Ablation Study
IV-D1 Effect of Unsupervised Training
@@ IV-D2 Effect of Changes on BERT and CSI-BERT
To demonstrate the effectiveness of the modifications we made on BERT [22], we first compare the recovery performance of the original BERT (where we simply use a linear layer as the embedding and output layer to support the CSI format) with our modified version, as shown in Fig. 7. We observe that the amplitude spectrum of the original BERT appears smooth across all subcarriers, indicating that it tends to map all tokens within each subcarrier to similar value. Although BERT can also achieve a relatively low MSE of 5.26 in WiGesture dataset, it fails to capture any valuable information.
@@
We then evaluate the efficacy of the introduced time embedding layer and the ARL for capturing temporal and spatial information, respectively, in comparison to CSI-BERT. The results of the ablation study for CSI classification and CSI prediction are shown in Fig. 6 and Table VIII, respectively. For the CSI classification task, we observed that the time embedding layer and ARL led to significant improvements in the WiFall and WiCount datasets, but not in the WiGesture dataset. This discrepancy may be related to the difficulty of the tasks: in the WiGesture dataset, the proposed method achieved nearly 100% accuracy even without pre-training, ARL, or the time embedding layer. In contrast, the proposed modules demonstrate greater benefits for the more complex tasks in the WiFall and WiCount datasets. For the CSI prediction task, both ARL and the time embedding layer significantly reduced prediction errors across all datasets. These findings illustrate that the two modules effectively help the model capture the temporal and spatial relationships within the data.
@@
(a) BERT[22] (sample set 1)
(b) CSI-BERT[11] (sample set 1)
@@
(d) BERT [22] (sample set 2)
(e) CSI-BERT[11] (sample set 2)
@@
(a) BERT[22] (2020-th packet)
(b) CSI-BERT[11] (2020-th packet)
@@
(d) BERT [22] (8080-th packet)
(e) CSI-BERT [11] (8080-th packet)
@@ V Conclusion
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with Texas A&M and University of California-Davis (Sept. 2016-Mar. 2018). After being a visiting research scholar (Apr. 2018 – June 2019) at Shenzhen Research Institute of Big Data, Shenzhen, China, he has been a research scientist since June 2019. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Xiaoyang Li (Member, IEEE) received the B.Eng. degree from the Southern University of Science and Technology (SUSTech) in 2016 and the Ph.D. degree from The University of Hong Kong in 2020. From 2020 to 2022, he was a Presidential Distinguished Research Fellow at SUSTech. He is currently an Associate Research Fellow with the Shenzhen Research Institute of Big Data, and an Adjunct Assistant Professor with the Chinese University of Hong Kong, Shenzhen. He is a recipient of the Youth Talent Support Program by China Communication Society, Forbes China 30 under 30, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications & Sensing, and the Exemplary Reviewer of IEEE Wireless Communications Letters. He served as the Editor of Journal of Information and Intelligence, the Workshop/Session Chairs of IEEE ICASSP/ICC/WCNC/PIMRC/MIIS. His research interests include integrated sensing-communication-computation, edge learning, and network optimization.

### v4

@@ IV-C2 CSI Prediction Task
Given the scarcity of real CSI datasets in wireless communications, we adopt the widely used approach of leveraging wireless sensing datasets for performance evaluation. For the prediction task, we provide the first 0.8 seconds of CSI in each sample and train the model to predict the subsequent 0.2 seconds of CSI. We compare the performance of our CSI-BERT2 model against several other CSI prediction models, including popular time series models like LSTM [LSTM], RNN [RNN], GRU [GRU], and Mamba [Mamba], which are also widely used in CSI prediction [LSTM_CSI1, LSTM_CSI2, LSTM_CSI3, LSTM_CSI4, RNN_CSI, GRU_CSI]. We also compare our model with some CSI prediction networks like OCEAN [OCEAN], CV-3DCNN [3DCNN], and ConvLSTM [ConvLSTM]. The prediction error and time cost (using GPU) are shown in Table VI. It can be seen that our CSI-BERT2 model significantly outperforms all other models. Particularly in the WiCount dataset, where the task is more challenging, as each gesture or action has a relatively irregular pattern, other methods exhibit a substantial decrease in performance, while CSI-BERT2 is influenced only minimally. Although the inference time of CSI-BERT2 is longer than that of the other models, its total inference time for the entire testing set across different datasets remains under 0.5 seconds, with a sample volume of 500 for each dataset using a sliding window. Consequently, the inference time for a single sample is within 1 ms, which is acceptable in practice. Additionally, we provide the Floating Point Operations (FLOPs) of different methods for a single input in the table. Our CSI-BERT2 model requires higher FLOPs, but it only reaches around 0.5G, making it suitable for real-time applications in edge devices [wang2025revisiting, wang2025ultra].
@@
Configuration CommPre WiGesture [CSI-BERT]
@@
LSTM [LSTM, LSTM_CSI1, LSTM_CSI2, LSTM_CSI3, LSTM_CSI4] 133K 0.0134 5.1107 0.4849 0.5722 0.03 12.3254 0.2397 0.3967 0.05
RNN [RNN, RNN_CSI] 33K 0.0033 4.6164 0.4607 0.5385 0.03 19.4708 0.2877 0.4063 0.04
GRU [GRU, GRU_CSI] 100K 0.0100 5.4009 0.5128 0.6367 0.04 19.7180 0.2922 0.4243 0.04
Mamba [Mamba] 5M 0.5062 5.1286 0.4430 0.4838 0.23 12.3281 0.2392 0.3277 0.24
OCEAN [OCEAN] 126K 0.0126 5.6013 0.4863 0.5576 0.05 19.6257 0.2925 0.4231 0.05
CV-3DCNN [3DCNN] 19K 0.1033 4.5276 0.4794 0.6810 0.04 11.3017 0.2267 0.3044 0.04
ConvLSTM [ConvLSTM] 152K 0.1136 5.0015 0.4898 0.6031 0.04 19.7038 0.2921 0.4242 0.04
Configuration WiFall[KNN-MMD] WiCount
@@
LSTM [LSTM, LSTM_CSI1, LSTM_CSI2, LSTM_CSI3, LSTM_CSI4] 133K 0.0134 7.1495 0.1624 0.1882 0.04 32.3377 0.2547 0.3528 0.05
RNN [RNN, RNN_CSI] 33K 0.0033 16.9083 0.2424 0.2988 0.04 32.3670 0.2548 0.3534 0.03
GRU [GRU, GRU_CSI] 100K 0.0100 16.5353 0.2395 0.2963 0.07 39.8108 0.2541 0.3556 0.02
Mamba [Mamba] 5M 0.5062 6.4666 0.1532 0.1756 0.12 39.9170 0.2566 0.3524 0.11
OCEAN [OCEAN] 126K 0.0126 16.8825 0.2423 0.2978 0.03 39.7917 0.2542 0.3548 0.02
CV-3DCNN [3DCNN] 19K 0.1033 8.2616 0.1713 0.1981 0.03 42.2662 0.2631 0.3560 0.02
ConvLSTM [ConvLSTM] 152K 0.1136 16.8935 0.2429 0.2983 0.03 39.7709 0.2537 0.3552 0.02
@@ IV-C3 CSI Classification Task
The robust and powerful capabilities of our CSI-BERT2 model in CSI classification tasks are demonstrated in Table V, where it achieves the best performance across all scenarios. Notably, the original CSI-BERT fails to outperform models such as Chen et al. [fall] and WiGrumt [WiGRUNT] in many scenarios. However, our CSI-BERT2 shows substantial improvement due to the enhanced model structure, thanks to the ARL capturing the relationships and importance among subcarriers, and the MLP-based time embedding layer that encodes time information more effectively. Moreover, aided by the time embedding mechanism, CSI-BERT2 can be effectively applied in scenarios where the training data includes samples with varying sampling rates, or where the sampling rate of the testing set differs from that of the training set. This is a practical feature, especially in federated learning scenarios, where we cannot guarantee that all clients have the same sampling rate. In Table VII, we compare the performance of different models where the training and testing sets both have 100Hz and 50Hz CSI, as well as where the training set only has 100Hz CSI but the testing set has 50Hz CSI. For fairness, we provide the CSI processed by linear interpolation for the other non-CSI-BERT models. It can be seen that other models are significantly influenced by the different sampling rates, with some even failing to function properly in particular scenarios. In contrast, CSI-BERT2 is almost unaffected by any impact, thanks to the help of time embedding.
@@
Data Training Set: 100Hz+50Hz; Testing Set: 100Hz+50Hz Model MLP CNN RNN [RNN] LSTM [LSTM] Chen et al. [fall] WiGRUNT [WiGRUNT] CSI-BERT [CSI-BERT] CSI-BERT2 Gesture Recognition 16.51% 14.54% 15.13% 17.37% 16.59% 74.47% 64.61% 97.04% Person Identification 13.47% 15.52% 13.41% 13.47% 13.72% 81.25% 70.83% 99.54% Action Recognition 70.97% 67.67% 70.02% 59.63% 75.18% 66.78% 78.18% 88.35% Fall Detection 81.10% 75.80% 83.92% 84.83% 85.98% 80.72% 92.98% 93.64% People Number Estimation 84.03% 46.82% 81.25% 82.00% 80.09% 76.13% 86.93% 92.54% Data Training Set: 100Hz; Testing Set: 50Hz Model MLP CNN RNN [RNN] LSTM [LSTM] Chen et al. [fall] WiGRUNT [WiGRUNT] CSI-BERT [CSI-BERT] CSI-BERT2 Gesture Recognition 69.79% 20.38% 36.25% 27.15% 74.89% 71.37% 79.96% 97.81% Person Identification 87.29% 11.85% 82.29% 22.32% 87.22% 85.24% 94.44% 99.38% Action Recognition 68.97% 51.96% 68.07% 60.26% 76.56% 73.21% 84.56% 88.53% Fall Detection 80.83% 76.13% 80.17% 84.38% 78.61% 77.08% 94.19% 94.32% People Number Estimation 82.22% 77.03% 84.43% 42.22% 68.89% 71.11% 89.10% 94.77%
@@ IV-D Ablation Study
IV-D1 Effect of Unsupervised Pre-Training
@@ IV-D2 Effect of Changes on BERT and CSI-BERT
To demonstrate the effectiveness of the modifications we made on BERT [BERT], we first compare the recovery performance of the original BERT (where we simply use a linear layer as the embedding and output layer to support the CSI format) with our modified version, as shown in Fig. 7. We observe that the amplitude spectrum of the original BERT appears smooth across all subcarriers, indicating that it tends to map all tokens within each subcarrier to similar value. Although BERT can also achieve a relatively low MSE of 5.26 in WiGesture dataset, it fails to capture any valuable information.
@@
We then evaluate the efficacy of the introduced time embedding layer and the ARL for capturing temporal and spatial information, respectively, in comparison to CSI-BERT. The results of the ablation study for CSI classification and CSI prediction are shown in Fig. 6 and Table VIII, respectively. For the CSI classification task, we observed that the time embedding layer and ARL led to significant improvements in the WiFall and WiCount datasets, but not in the WiGesture dataset. This discrepancy may be related to the difficulty of the tasks: in the WiGesture dataset, the proposed method achieved nearly 100% accuracy even without pre-training, ARL, or the time embedding layer. In contrast, the proposed modules demonstrate greater benefits for the more complex tasks in the WiFall and WiCount datasets, , where the ARL-based spatial embedding and MLP-based time embedding resulted in improvements of approximately 13.90% and 12.79%, respectively. For the CSI prediction task, both the ARL and the time embedding layer significantly reduced prediction errors across all datasets, resulting in approximately 15.47% and 10.95% reduction in MSE, respectively, among the four datasets. These findings illustrate that the two modules effectively help the model capture the temporal and spatial relationships within the data.
@@
(a) BERT[BERT] (sample set 1)
(b) CSI-BERT[CSI-BERT] (sample set 1)
@@
(d) BERT [BERT] (sample set 2)
(e) CSI-BERT[CSI-BERT] (sample set 2)
@@
(a) BERT[BERT] (2020-th packet)
(b) CSI-BERT[CSI-BERT] (2020-th packet)
@@
(d) BERT [BERT] (8080-th packet)
(e) CSI-BERT [CSI-BERT] (8080-th packet)
@@ V Conclusion
[Uncaptioned image] Hang Li received the B.E. and M.S. degrees from Beihang University, Beijing, China, in 2008 and 2011, respectively, and the Ph.D. degree from Texas A&M University, College Station, TX, USA, in 2016. He was a postdoctoral research associate with both Texas A&M University and University of California-Davis (Sept. 2016-Mar. 2018). He was a visiting research scholar (Apr. 2018 – June 2019) and a research scientist (June 2019 – May 2025) at Shenzhen Research Institute of Big Data, Shenzhen, China. His current research interests include wireless networks, Internet of things, stochastic optimization, and applications of machine learning. He is recognized as Overseas High-Caliber Personnel (Level C) at Shenzhen in 2020.
[Uncaptioned image] Xiaoyang Li (Member, IEEE) is currently an Assistant Professor with Southern University of Science and Technology. He received the Ph.D. degree from The University of Hong Kong. His research interests include integrated sensing-communication-computation and edge learning. He is a recipient of Young Elite Scientists Sponsorship Program by CAST, Forbes China 30 under 30, Young Elite of G20, Overseas Youth Talent in Guangdong, Overseas High-caliber Personnel in Shenzhen, Outstanding Research Fellow in Shenzhen, the Best Paper Award of IEEE 4th International Symposium on Joint Communications and Sensing, the Exemplary Reviewers of IEEE Wireless Communications Letters and Journal of Information and Intelligence (JII). He has served as the Editor of JII, and the Workshop Chairs of IEEE ICASSP/WCNC/PIMRC/MIIS/MediCom.